# Dataset Distillation-based Hybrid Federated Learning on Non-IID Data

### v1

@@ V-C Communication Cost
CIFAR10 ($C=1$) 54.4% 61.9% 60.6% 58.7% 61.8% 64.3% 61.3%
CIFAR10 ($C=2$) 70.8% 69.7% 70.8% 69.1% 68.9% 70.1% 68.2%
CIFAR10 ($C=10$) 73.5% 74.5% 74.4% 73.1% 72.4% 72.2% 71.3%
@@
V-D Impact of $K$ in K-Means Clustering
In this part, we evaluate the impact of $K$ in the K-Means homogeneous client clustering on the performance of the proposed HFLDD. Table IV shows the test accuracies of HFLDD on the MNIST and CIFAR10 datasets under different values of $K$. When $C=1$, on the CIFAR10 dataset, the test accuracy is the highest when $K=10$, and it shows a noticeable decline as $K$ decreases or increases. However, on the MNIST dataset, the test accuracy remains relatively unchanged as $K$ varies. This is because the MNIST dataset is relatively simple, making its data patterns and features easy to capture and understand. In this scenario, even if an inappropriate value of $K$ is selected, HFLDD still can easily learn the features in the MNIST dataset. When $C=2$ and $C=10$, the heterogeneity is mitigated, and there is partial overlap in label classes among different clients, leading to a certain degree of overlap in the feature space. The test accuracy on the two datasets remains relatively stable as the value of $K$ changes. From this, we observe that when data heterogeneity is not particularly severe, HFLDD exhibits robustness to the value of $K$.
VI Discussion
@@
It is important to note that the actual scenario differs from the example provided above, the data heterogeneity among clients’ data types is highly complex. After sampling and reorganizing homogeneous clusters, it is highly possible to have a limited number of clients in some heterogeneous clusters. There may even be many isolated nodes considered as a single heterogeneous cluster. This leads to uneven number of clients among clusters. In decentralized settings, if these uneven clusters are linked, as referenced in [30], to form a ring-shaped topology for collaborative training, it will lead to a significant decrease in convergence speed and test accuracy. The most intuitive approach to address this issue is to directly remove the heterogeneous clusters with isolated or a very small number of clients from training. However, this often results in worse training performance, failing to fully utilize the data in each client.
@@
For better understanding of this issue in HFL, we let each client hold two different classes of data. Due to the overlap of data classes among clients, it is likely to form homogeneous clusters as illustrated in Fig. 7a. In this case, through sampling and reorganization, the constructed heterogeneous clusters are with unbalanced client numbers, as shown in Fig. 7b. After transmitting the distilled data, heterogeneous clusters with fewer clients cannot obtain a dataset whose distribution approximates to the global distribution.
However, unlike the decentralized learning in [30], in HFL, the impact of these heterogeneous clusters with very few clients or even just isolated nodes on the performance of model training is relatively small. This is mainly because the presence of a central server makes data exchange and collaboration more flexible. Even if some clusters have only a very few clients, the server can still integrate and utilize their information, thus making full use of all data resources without excluding these imbalanced clusters.
VII Conclusion
@@
In this paper, we propose a communication-efficient HFL framework that addresses the challenge of distributed learning with Non-IID data by leveraging dataset distillation. Our approach constructs heterogeneous clusters based on label knowledge from all clients. Within each heterogeneous cluster, members transmit distilled data to a cluster header. This process aims to generate approximately IID datasets among clusters. Extensive experiments demonstrate that our proposed HFLDD outperforms the considered baseline methods, and can significantly reduce the impact of Non-IID data on training performance with a much smaller communication cost.

### v2

@@ VI-E Impact of Network Architecture
CIFAR10 ($C=1$) 64.1% 71.6% 70.3% 68.4% 71.5% 74.0% 71.0%
CIFAR10 ($C=2$) 79.0% 78.1% 79.0% 77.5% 77.3% 78.5% 76.6%
CIFAR10 ($C=10$) 78.5% 79.4% 79.3% 79.7% 79.1% 78.8% 77.9%
CINIC10 ($C=1$) 51.8% 51.5% 55.3% 54.8% 54.4% 56.5% 55.2%
CINIC10 ($C=2$) 57.9% 58.1% 58.5% 57.8% 57.5% 59.5% 59.7%
CINIC10 ($C=10$) 62.8% 63.5% 63.7% 62.8% 62.9% 62.6% 62.0%
VI-F Impact of $K$ in K-Means clustering
In this part, we evaluate the impact of $K$ in the homogeneous clustering of K-Means clients on the performance of the proposed HFLDD. Table VI shows the test accuracies of HFLDD in the MNIST, CIFAR10 and CINIC10 datasets with different values of $K$. When $C=1$, on the CIFAR10 and CINIC10 dataset, the test accuracy is the highest when $K=10$, and shows a noticeable decline as $K$ decreases or increases. However, in the MNIST dataset, the test accuracy remains relatively unchanged as $K$ varies. This is because the MNIST dataset is relatively simple, making its data patterns and features easy to capture and understand. In this scenario, even if an inappropriate value of $K$ is selected, HFLDD still can easily learn the features in the MNIST dataset.
When $C=2$ and $C=10$, heterogeneity is mitigated and there is partial overlap in label classes between different clients, leading to a certain degree of overlap in the feature space. The test accuracy on the three datasets remains relatively stable as the value of $K$ changes. From this, we observe that when the heterogeneity of the data is not particularly severe, HFLDD exhibits robustness to the value of $K$.
In addition to model accuracy, we further examine the effect of $K$ on the convergence rate. As illustrated in Fig. 8. From the figure, we observe that the number of clusters has a relatively small effect on the convergence speed.
(a) $C=1$
(b) $C=2$
(c) $C=10$
Figure 8: Learning curves of the global model under different homogeneous cluster numbers across varying data heterogeneity on CIFAR10.
@@ VI-F Impact of $K$ in K-Means clustering
VII Discussion
@@
It is important to note that the actual scenario differs from the example provided above, the data heterogeneity among clients’ data types is highly complex. After sampling and reorganizing homogeneous clusters, it is highly possible to have a limited number of clients in some heterogeneous clusters. There may even be many isolated nodes considered as a single heterogeneous cluster. This leads to an uneven number of clients among clusters. In decentralized settings, if these uneven clusters are linked, as in Ref. [32], to form a ring-shaped topology for collaborative training, it will lead to a significant decrease in convergence speed and test accuracy. The most intuitive approach to address this issue is to directly remove heterogeneous clusters with isolated or a very small number of clients from training. However, this often results in worse training performance, which is unable to fully utilize the data in each client.
@@
For better understanding of this issue in HFL, we let each client hold two different classes of data. Due to the overlap of data classes between clients, it is likely to form homogeneous clusters as illustrated in Fig. 9a. In this case, through sampling and reorganization, the constructed heterogeneous clusters have unbalanced client numbers, as shown in Fig. 9b. After transmitting the distilled data, heterogeneous clusters with fewer clients cannot obtain a dataset whose distribution approximates the global distribution.
Although some heterogeneous clusters may have uneven client sizes and lack complete label coverage, HFLDD is designed to robustly tolerate such imperfections. This robustness is mainly attributed to three factors. First, it promotes statistical diversity by sampling clients from different homogeneous groups. Although full label coverage within each heterogeneous cluster is not guaranteed, this strategy encourages the inclusion of diverse label types. As a result, many clusters exhibit approximately balanced label distributions in practice, which is both realistic and sufficient for achieving robust training, particularly under severe non-IID conditions. Second, the presence of a central server makes data exchange and collaboration more flexible. Each client participates in at least one heterogeneous cluster, and through multiround global aggregation, the distilled knowledge from all clients is eventually integrated into the global model. This ensures that no client is excluded from the learning process. Third, while uneven cluster sizes result in different numbers of distilled data across cluster heads, our experiments show that such an imbalance has a negligible effect on the final model performance. Therefore, HFLDD maintains stable performance despite inevitable imperfections in cluster formation, making it particularly suitable for real-world implementation.
VIII Conclusion
@@
In this paper, we propose a communication-efficient HFL framework that addresses the challenge of distributed learning with non-IID data by leveraging dataset distillation. Our approach constructs heterogeneous clusters based on the knowledge of the labels of all clients. Within each heterogeneous cluster, members transmit distilled data to a cluster head. This process aims to generate approximately IID datasets among clusters. Extensive experiments demonstrate that our proposed HFLDD outperforms the considered baseline methods and can significantly reduce the impact of non-IID data on training performance with a much smaller communication cost.